OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 1024 and a3 = 64?

OpenStudy (anonymous):

Hey - so in a geometric sequence, you're multiplying each term by the same thing to get to the next thing a1 = 1024 a2 = ? a3 = 64 a4...a8

OpenStudy (anonymous):

so what's 64's relation to 1024? How do you get to 64 in two steps?

OpenStudy (anonymous):

I am trying to figure that out... 1024/4=256/4=64...

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

so then keep dividing by 4

OpenStudy (anonymous):

a3 = 64 a4 = 64/4 = 16 a5 = 16/4 ... and so on

OpenStudy (anonymous):

64/4=16/4=4 a4 4/4=1/4=.25 a5 .25/4=.0625/4=.015625 a6

OpenStudy (anonymous):

Wait use mathematics here..

OpenStudy (anonymous):

...

OpenStudy (anonymous):

watch out on a5 ... and keep them in fraction form. will make it easier. a3 = 64 a4 = 64/4 = 16 a5 = 16/4 = ...

OpenStudy (anonymous):

\[a = 1024\] \[a_3 = 64\] \[\large 64 = a \cdot r^{3-1} \implies (1024) \cdot (r^2)\] Find r from it..

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

\[\large r^2 = \frac{64}{1024} \implies r^2 = \frac{1}{16} \implies r = \frac{1}{4}\]

OpenStudy (anonymous):

Now you have to find 8th term: \[\large a_8 = a \cdot (r)^{8-1} \implies a \cdot (r)^7\] Can you find now ??

OpenStudy (anonymous):

.0625?

OpenStudy (anonymous):

\[\large a_8 = (1024) \times \frac{1}{4^7} \implies \color{green}{ 0.0625}\] Yes you are right..

OpenStudy (anonymous):

Thank You Both!